Fermi National Laboratory


Accelerator Update for August 15, 2005 - August 17, 2005

Monday, August 15
The day shift began with Operations monitoring store 4332 and the Recycler stash, with the Antiproton Source (Pbar) in access to repair its lithium lens, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE, NuMI, and the Meson MCenter and MTest experiments taking beam.

Pbar experts accessed their vault in the AP1 transfer line, verified that the problem was with the lens, and began repairs at 10:19 AM.

A Main Injector (MI) expert began a study at 10:19 AM.

Beginning at 3 PM, Booster experts replaced a power supply (B:MV2) and Preaccelerator experts accessed the I- Source and replaced its generator brushes. Both of these jobs were completed by 3:44 PM and beam resumed.

A NuMI expert began a bunch rotation study at 3:51 PM.

Pbar experts reported at 9:30 PM that they had completed pulse testing the new lithium lens and were closing up the vault.

At 9:52 PM, the Tevatron (TeV) suffered a sector F1 quench. The store was lost. Beam was inadvertently sent from Pbar to an F sector magnet, which caused the quench; it also caused F sector vacuum problems.

At 10:10 PM, the Fermilab Tevatron motion sensors at CDF and D0 detected the earthquake that hit Honshu, Japan at 2146 CDT.

Operators prepared the CDF and D0 collision halls for access at 23:03 PM.

Tuesday, August 16
The midnight shift began with the TeV recovering from a quench and with vacuum techs in F sector, and with Operations monitoring the Recycler stash and the Pbar stack

CDF and D0 were both out of access by 12:04 AM.

Cryo system technicians reported at 12:41 AM that they had finished their work on the TeV sector F2 system and were moving to F4 to work on that expander.

At 1:52 AM, Operations reported that an access would be needed into F sector to look for a vacuum leak.

Vacuum techs reported at 6:27 AM that they had found a bad O-ring in F sector and were beginning the repairs.

Operators prepared Booster for an access at 7:32 AM so that RF techs could repair BRF9.

D0 experimenters accessed their collision hall from 9:21 AM to 10:21 AM.

Operations put the TeV through a dry squeeze at 10:35 AM.

Booster RF experts reported at 11 AM that they found a bad cable between BRF9ís modulator and power amplifier. They bypassed the station while they finished their work out of the tunnel.

Operations established beam to NuMI at 11:53 AM.

Operations injected some beam into the TeV at 12:15 PM for a tune up.

Pbar resumed stacking at 12:26 PM.

Booster RF experts finished their work on BRF9 and tied it back into the system at 2:56 PM. Pbar, NuMI, MiniBooNE, and Switchyard (SY) all resumed taking beam at 3:09 PM.

A MI expert began a SY fast spill study at 3:24 PM.

At 4:22 PM, BRF16 tripped off twice in a couple of minutes. An expert investigated and had the station back up twelve minutes later.

MCenter and MTest resumed taking beam at 4:32 PM.

Operations put the TeV through a wet squeeze at 8:17 PM.

Operations put the TeV into shot setup at 8:45 PM.

Operations established combination store 4336 at 11:12 PM with an initial luminosity of 76.29E30.

At 11:20 PM, all the MI RF stations tripped off due to a frequency problem. It did it again nineteen minutes later. The RF system reset both times.

Operations established beam to MCenter and MTest at 11:21 PM.

Pbar resumed stacking at 11:22 PM.

Wednesday, August 17
The midnight shift began with Operations monitoring store 4336, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 12:51 AM, a Linac RF station (LRF3) tripped off for the first of four times in two hours. It reset each time.

A Recycler expert began a study at 1:28 AM.

Operations put the Recycler into shot setup at 3:41 AM to stash antiprotons.

At 5:25 AM, Operations reported that the last antiproton transfer didnít make it into the Recycler, nor did it circulate in MI.

Pbar resumed stacking at 5:42 AM.

MCenter resumed taking beam at 5:45 AM.

The Plans for Wednesday
The plan for today is to store, stack, and stash.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/17/2005   email Fermilab